# RF-DETR: Neural Architecture Search for Real-Time Detection Transformers

### v1

@@ 4 Experiments
We evaluate RF-DETR on COCO and RF100-VL and demonstrate that our approach achieves state-of-the-art accuracy among all real-time methods. In addition, we identify inconsistencies in standard benchmarking protocols and present a simple standardized procedure to improve reproducibility. Following LW-DETR (chen2024lw), we group models of similar latency into the same size bucket rather than grouping based on parameter count.
@@
Table 1: Standardizing Latency Evaluation. Variance in latency measurements can be largely attributed to power throttling and GPU overheating. We mitigate this issue by buffering for 200ms between forward passes. Notably, this benchmarking approach is not designed to measure sustained throughput, but rather ensures reproducible latency measurements. We are unable to reproduce YOLOv8 and YOLOv11’s mAP results in TensorRT, likely because these models evaluate with multi-class NMS but only use single-class NMS in inference. We use the standard NMS-tuned confidence threshold of 0.01. YOLOv8 and YOLOv11 performance degrades further when quantizied from FP32 to FP16, reaffirming that all models should report latency and accuracy using the same model artifact. Notably, naively quantizing D-FINE to FP16 reduces performance to 0.5 AP. We fix this issue by changing the authors’ export code to use ONNX opset 17. See Appendix A for more details.
@@
Standardizing Latency Benchmarking. Despite its maturity, benchmarking object detectors remains inconsistent across prior work. For example, YOLO-based models often omit non-maximal suppression (NMS) when computing latency, leading to unfair comparisons with end-to-end detectors. Additionally, YOLO-based segmentation models measure the latency of generating prototype predictions instead of directly usable per-object masks (yolov11), leading to biased runtime measurements. Further, D-FINE’s reported latency evaluation of LW-DETR is 25% faster than reported by chen2024lwdetr. We observe that such differences can be attributed to detectable power throttling events, particularly when the GPU overheats (cf. Table 1). In contrast, simply pausing for 200ms between consecutive forward passes largely mitigates power throttling, yielding more stable latency measurements. Lastly, we find that prior work often reports latency using FP16 quantized models, but evaluates accuracy with FP32 models. However, naive quantization can significantly degrade performance (in some cases dropping performance to near 0 AP). To ensure fair comparison, we advocate reporting accuracy and latency with the same model artifact. We release our stand-alone benchmarking tool on GitHub.
@@
YOLOv8 (yolov8) N 3.2M 8.7 2.1 35.2 49.2 38.3 15.8 38.8 51.3
YOLOv11 (yolov11) N 2.6M 6.5 2.2 37.1 51.6 40.4 17.3 40.7 55.6
YOLOv8 (yolov8) S 11.2M 28.6 2.9 42.4 57.6 46.0 22.2 47.1 59.6
YOLOv11 (yolov11) S 9.4M 21.5 3.2 44.1 59.3 47.9 26.1 48.5 62.6
YOLOv8 (yolov8) M 25.9M 78.9 5.4 47.3 62.5 51.5 27.5 52.9 65.1
YOLOv11 (yolov11) M 20.1M 68.0 5.1 48.3 63.6 52.5 29.1 53.8 66.3
@@
LW-DETR (chen2024lw) T 12.1M 21.4 1.9 42.9 60.7 45.9 22.7 47.3 60.0
D-FINE (peng2024dfine) N 3.8M 7.3 2.1 42.7 60.2 45.4 22.9 46.6 62.1
@@
LW-DETR (chen2024lw) S 14.6M 31.8 2.6 48.0 66.8 51.6 26.7 52.5 65.6
D-FINE (peng2024dfine) S 10.2M 25.2 3.5 50.6 67.6 55.0 32.6 54.6 66.6
@@
RT-DETR (zhao2024rtdetr) R18 36.0M 100.0 4.4 49.0 66.6 53.3 32.8 52.1 65.0
LW-DETR (chen2024lw) M 28.2M 83.9 4.4 52.6 72.0 56.6 32.5 57.6 70.5
D-FINE (peng2024dfine) M 19.2M 56.6 5.4 55.0 72.6 59.7 37.6 59.4 71.7
@@
Evaluating RF-DETR and RF-DETR-Seg on COCO. COCO (lin2014coco) is a flagship benchmark for object detection and instance segmentation. In Table 2, we compare RF-DETR with leading real-time and open-vocabulary detectors. RF-DETR (nano) beats both D-FINE (nano) and LW-DETR (nano) by more than 5 AP. We see similar trends for small and medium sizes as well. Notably, RF-DETR also significantly outperforms YOLOv8 and YOLOv11. RF-DETR (nano) matches the performance of YOLOv8 and YOLOv11 (medium). We use mmdetection’s implementation of GroundingDINO and include their reported AP since they do not release a model artifact for GroundingDINO fine-tuned on COCO. We benchmark mmGroundingDINO’s parameter count, GFLOPS, and latency using the released open-vocabulary model. In Table 3, we compare RF-DETR-Seg with real-time instance segmentation models. RF-DETR-Seg (nano) outperforms YOLOv8 and YOLOv11 at all sizes. Furthermore, RF-DETR-Seg (nano) beats FastInst by 5.4% while running almost ten times faster. Similarly, RF-DETR (x-large) surpasses GroundingDINO (tiny), and RF-DETR-Seg (large) outperforms MaskDINO (R50), at a fraction of their runtime.
Table 3: COCO Instance Segmentation Evaluation. We compare RF-DETR with popular real-time instance segmentation methods on COCO. Notably, RF-DETR (nano) outperforms all reported YOLOv8 and YOLOv11 model sizes. Further RF-DETR (nano) outperforms FastInst by 4.4%, while running nearly ten times faster. RF-DETR(medium) approaches the performance on MaskDINO at a fraction of the runtime. We denote models that do not support TensorRT execution with a star, and instead report PyTorch latency results. Our latencies for YOLOs also include the conversion of protos into masks, which are not typically included in prior benchmarks but nonetheless contribute meaningfully to practical latency. See Appendix E for L, XL, and Max variants of RF-DETR-Seg on COCO.
@@
YOLOv8 (yolov8) N 3.4M 12.6 3.5 28.3 45.6 29.8 9.3 31.3 44.3
YOLOv11 (yolov11) N 2.9M 10.4 3.6 30.0 47.8 31.5 10.0 33.4 47.7
YOLOv8 (yolov8) S 11.8M 42.6 4.2 34.0 53.8 36.0 13.6 38.5 52.2
YOLOv11 (yolov11) S 10.1M 35.5 4.6 35.0 55.4 37.1 15.3 39.7 53.9
YOLOv8 (yolov8) M 27.3M 110.2 7.0 37.3 58.2 39.9 16.7 43.0 56.1
YOLOv11 (yolov11) M 22.4M 123.3 6.9 38.5 60.0 40.9 18.0 44.3 57.6
@@
FastInst (he2023fastinst) R50 29.7M 99.7 39.6 34.9 56.0 36.2 13.3 38.0 56.8
MaskDINO (li2023maskdino) R50 52.1M 586 242 46.3 69.0 50.7 26.1 49.3 66.1
@@
Evaluating RF-DETR on RF100-VL. RF100-VL is a challenging detection benchmark composed of 100 diverse datasets. We report latencies, FLOPs, and accuracy averaged over all 100 datasets in Table 4. Our results show that RF-DETR (2x-large) outperforms GroundingDINO and LLMDet while requiring only a fraction of their runtime. Interestingly, RT-DETR outperforms D-FINE (which is built on RT-DETR) at mAP50, indicating that D-FINE’s hyperparameters are potentially overoptimized for COCO. We note that RF-DETR benefits from scaling to larger backbone sizes (Appendix E). In contrast, YOLOv8 and YOLOv11 consistently underperform DETR-based detectors, and scaling these model families to larger sizes does not improve their performance on RF100-VL.
@@
YOLOv8 (yolov8) N 3.2M 8.7 2.6 55.0 81.1 59.5 4.8 44.1 48.0
YOLOv11 (yolov11) N 2.6M 6.5 3.0 55.5 81.3 60.3 4.7 44.4 49.2
YOLOv8 (yolov8) S 11.2M 28.6 3.1 56.3 82.0 60.9 6.1 45.6 48.6
YOLOv11 (yolov11) S 9.4M 21.5 3.3 56.4 82.5 61.3 6.5 45.5 48.5
YOLOv8 (yolov8) M 25.9M 78.9 5.4 56.5 82.3 60.9 6.4 45.7 48.6
YOLOv11 (yolov11) M 20.1M 68.0 5.1 57.0 82.5 61.9 7.3 46.1 48.6
@@
GroundingDINO (liu2023grounding) T 173.0M 1008.3 309.9 62.3 88.8 67.8 39.2 57.7 69.5
LLMDet (fu2025llmdet) T 173.0M 1008.3 308.4 62.3 88.3 67.8 39.1 57.6 70.3
@@
LW-DETR (chen2024lw) N 12.1M 21.4 1.9 57.1 84.7 61.5 31.2 51.8 65.8
D-FINE (peng2024dfine) N 3.8M 7.3 2.0 58.2 84.4 62.5 32.4 52.9 65.8
RF-DETR (Ours) N 31.2M 34.5 2.5 57.6 84.9 62.1 30.7 52.2 66.8
RF-DETR w/ Fine-Tuning (Ours) N 31.2M 34.5 2.5 58.7 85.6 63.5 32.4 52.7 67.0
LW-DETR (chen2024lw) S 14.6M 31.8 2.6 57.4 85.0 62.0 32.1 52.1 65.8
D-FINE (peng2024dfine) S 10.2M 25.2 3.5 60.3 85.3 65.4 36.6 56.0 68.4
RF-DETR (Ours) S 33.5M 62.4 3.7 60.7 87.0 66.0 35.4 55.4 69.6
RF-DETR w/ Fine-Tuning (Ours) S 33.5M 62.4 3.7 61.0 87.2 66.4 35.3 55.9 69.8
RT-DETR (zhao2024rtdetr) M 36.0M 100.0 4.3 59.6 85.7 64.6 36.4 54.6 67.3
LW-DETR (chen2024lw) M 28.2M 83.9 4.3 59.8 86.8 64.9 34.0 54.4 68.9
D-FINE (peng2024dfine) M 19.2M 56.6 5.6 60.6 85.5 65.8 36.0 56.6 67.5
RF-DETR (Ours) M 33.5M 86.7 4.6 61.5 87.7 67.0 36.44 56.5 69.8
RF-DETR w/ Fine-Tuning (Ours) M 33.5M 86.7 4.6 61.9 87.9 67.3 36.4 56.6 70.1
@@
Table 5: Ablation on Neural Architecture Search. We ablate the impact of each “tunable knob” on accuracy and latency below. Using a gentler set of hyperparameters compared to LW-DETR (e.g. smaller batch size, lower learning rate, replacing batch norm with layer norm) reduces performance by 1%. However, we regain this lost performance by replacing LW-DETR’s CAEV2 backbone with DINOv2. Importantly, the lower learning rate and layer-norm allow us to better preserve DINOv2’s foundational knowledge and allows us to train with larger batch sizes, making weight-sharing NAS more effective. Counterintuitively, introducing weight sharing NAS to the training scheme improves performance of the base configuration even though patch size 14 isn’t in the NAS search space.
@@
Impact of Backbone Architecture and Pre-Training. We study the impact of different backbone architectures in RF-DETR. We find that DINOv2 achieves the best performance, outperforming CAEv2 by 2%. Interestingly, despite having fewer parameters than SigLIPv2, SAM2’s Hiera-S backbone is considerably slower. This is in contrast with the Hiera-S claim that it is meaningfully faster than equivalently performant ViTs. However, Hiera does not explore latency in the context of kernels such as Flash Attention, which are highly optimized in compilers such as TensorRT. Additionally, existing foundation model families typically do not release lightweight ViT variants such as ViT-S or ViT-T, making it difficult to repurpose such models for real-time applications.
Table 6: Ablation on Backbone. We ablate the impact of using different backbone architectures for RF-DETR below. We find that DINOv2 achieves the highest performance, outperforming CAEv2 by 2.4%. All models are pretrained with 60 epochs of Objects365 and the ’Gentler Hyperparameters’ setting. Note that SAM2 and SigLIPv2 perform poorly when evaluated in FP16. Therefore, we report FP16 TensorRT latency with FP32 ONNX accuracy for these two models as an upper bound on what their performance could be if optimized for FP16.
@@
Rethinking Standard Accuracy Benchmarking Practices. Following prior work, we report all COCO results on the validation set. However, relying solely on the validation for both model selection and evaluation can lead to overfitting. For example, D-FINE (which builds on RT-DETR) conducts an extensive hyperparameter sweep on COCO’s validation set and reports its best model. However, evaluating this configuration on RF100-VL shows that D-FINE underperforms RT-DETR on the test set. In contrast, our method achieves state-of-the-art performance among all real-time detectors on RF100-VL and COCO, demonstrating the robustness of our weight-sharing NAS. In addition to evaluating on COCO, we advocate that future detectors should also evaluate on datasets with public validation and test splits like RF100-VL.
Limitations. Despite controlling for power throttling and GPU overheating during inference, our latency measurements still have a variance of up to 0.1ms due to the non-deterministic behavior of TensorRT during compilation. Specifically, TensorRT can introduce power throttling, which in turn affects the resulting engine and leads to random fluctuations in latency. Although the measurement of a given TensorRT engine is generally consistent, recompiling the same ONNX artifact can produce different latency results. Therefore, we only report latencies with one digit of precision after the decimal place.

### v2

@@ 4 Experiments
We evaluate RF-DETR on COCO and RF100-VL and demonstrate that our approach achieves state-of-the-art accuracy among all real-time methods. In addition, we identify inconsistencies in standard benchmarking protocols and present a simple standardized procedure to improve reproducibility. Following LW-DETR (Chen et al., 2024a), we group models of similar latency into the same size bucket rather than grouping based on parameter count.
@@
Table 1: Standardizing Latency Evaluation. Variance in latency measurements can be largely attributed to power throttling and GPU overheating. We mitigate this issue by buffering for 200ms between forward passes. Notably, this benchmarking approach is not designed to measure sustained throughput, but rather ensures reproducible latency measurements. We are unable to reproduce YOLOv8 and YOLOv11’s mAP results in TensorRT, likely because these models evaluate with multi-class NMS but only use single-class NMS in inference. We use the standard NMS-tuned confidence threshold of 0.01. YOLOv8 and YOLOv11 performance degrades further when quantizied from FP32 to FP16, reaffirming that all models should report latency and accuracy using the same model artifact. Notably, naively quantizing D-FINE to FP16 reduces performance to 0.5 AP. We fix this issue by changing the authors’ export code to use ONNX opset 17 (Appendix A).
@@
Standardizing Latency Benchmarking. Despite its maturity, benchmarking object detectors remains inconsistent across prior work. For example, YOLO-based models often omit non-maximal suppression (NMS) when computing latency, leading to unfair comparisons with end-to-end detectors. Additionally, YOLO-based segmentation models measure the latency of generating prototype predictions instead of directly usable per-object masks (Jocher et al., 2024), leading to biased runtime measurements. Further, D-FINE’s reported latency evaluation of LW-DETR is 25% faster than reported by Chen et al. (2024b). We observe that such differences can be attributed to detectable power throttling events, particularly when the GPU overheats (Table 1). In contrast, simply pausing for 200ms between consecutive forward passes largely mitigates power throttling, yielding more stable latency measurements (Appendix K). Lastly, we find that prior work often reports latency using FP16 quantized models, but evaluates accuracy with FP32 models. However, naive quantization can significantly degrade performance (in some cases dropping performance to near 0 AP). To ensure fair comparison, we advocate for reporting accuracy and latency with the same model artifact. We release our stand-alone benchmarking tool on GitHub.
@@
YOLOv8 (Jocher et al., 2023) N 3.2M 8.7 2.1 35.2 49.2 38.3 15.8 38.8 51.3
YOLOv11 (Jocher et al., 2024) N 2.6M 6.5 2.2 37.1 51.6 40.4 17.3 40.7 55.6
YOLOv8 (Jocher et al., 2023) S 11.2M 28.6 2.9 42.4 57.6 46.0 22.2 47.1 59.6
YOLOv11 (Jocher et al., 2024) S 9.4M 21.5 3.2 44.1 59.3 47.9 26.1 48.5 62.6
YOLOv8 (Jocher et al., 2023) M 25.9M 78.9 5.4 47.3 62.5 51.5 27.5 52.9 65.1
YOLOv11 (Jocher et al., 2024) M 20.1M 68.0 5.1 48.3 63.6 52.5 29.1 53.8 66.3
@@
LW-DETR (Chen et al., 2024a) T 12.1M 21.4 1.9 42.9 60.7 45.9 22.7 47.3 60.0
D-FINE (Peng et al., 2024) N 3.8M 7.3 2.1 42.7 60.2 45.4 22.9 46.6 62.1
@@
LW-DETR (Chen et al., 2024a) S 14.6M 31.8 2.6 48.0 66.8 51.6 26.7 52.5 65.6
D-FINE (Peng et al., 2024) S 10.2M 25.2 3.5 50.6 67.6 55.0 32.6 54.6 66.6
@@
RT-DETR (Zhao et al., 2024) R18 36.0M 100.0 4.4 49.0 66.6 53.3 32.8 52.1 65.0
LW-DETR (Chen et al., 2024a) M 28.2M 83.9 4.4 52.6 72.0 56.6 32.5 57.6 70.5
D-FINE (Peng et al., 2024) M 19.2M 56.6 5.4 55.0 72.6 59.7 37.6 59.4 71.7
@@
Evaluating RF-DETR and RF-DETR-Seg on COCO. COCO (Lin et al., 2014) is a flagship benchmark for object detection and instance segmentation. In Table 2, we compare RF-DETR with leading real-time and open-vocabulary detectors. RF-DETR (nano) beats both D-FINE (nano) and LW-DETR (nano) by more than 5 AP. We see similar trends for small and medium sizes as well. Notably, RF-DETR also significantly outperforms YOLOv8 and YOLOv11. RF-DETR (nano) matches the performance of YOLOv8 and YOLOv11 (medium). We use mmdetection’s implementation of GroundingDINO and include their reported AP since they do not release a model artifact for GroundingDINO fine-tuned on COCO. We benchmark mmGroundingDINO’s parameter count, GFLOPS, and latency using the released open-vocabulary model. In Table 3, we compare RF-DETR-Seg with real-time instance segmentation models. RF-DETR-Seg (nano) outperforms YOLOv8 and YOLOv11 at all sizes. Furthermore, RF-DETR-Seg (nano) beats FastInst by 5.4% while running almost ten times faster. Similarly, RF-DETR (x-large) surpasses GroundingDINO (tiny), and RF-DETR-Seg (large) outperforms MaskDINO (R50), at a fraction of their runtime.
Table 3: COCO Instance Segmentation Evaluation. We compare RF-DETR with popular real-time instance segmentation methods on COCO. Notably, RF-DETR (nano) outperforms all reported YOLOv8 and YOLOv11 model sizes. Further RF-DETR (nano) outperforms FastInst by 5.4%, while running nearly ten times faster. RF-DETR (medium) approaches the performance on MaskDINO at a fraction of the runtime. We denote models that do not support TensorRT execution with a star, and instead report PyTorch latency results. Our latencies for YOLOs also include the conversion of protos into masks, which are not typically included in prior benchmarks but nonetheless contribute meaningfully to practical latency. See Appendix E for L, XL, and Max variants of RF-DETR-Seg on COCO.
@@
YOLOv8 (Jocher et al., 2023) N 3.4M 12.6 3.5 28.3 45.6 29.8 9.3 31.3 44.3
YOLOv11 (Jocher et al., 2024) N 2.9M 10.4 3.6 30.0 47.8 31.5 10.0 33.4 47.7
YOLOv8 (Jocher et al., 2023) S 11.8M 42.6 4.2 34.0 53.8 36.0 13.6 38.5 52.2
YOLOv11 (Jocher et al., 2024) S 10.1M 35.5 4.6 35.0 55.4 37.1 15.3 39.7 53.9
YOLOv8 (Jocher et al., 2023) M 27.3M 110.2 7.0 37.3 58.2 39.9 16.7 43.0 56.1
YOLOv11 (Jocher et al., 2024) M 22.4M 123.3 6.9 38.5 60.0 40.9 18.0 44.3 57.6
@@
FastInst (He et al., 2023) R50 29.7M 99.7 39.6 34.9 56.0 36.2 13.3 38.0 56.8
MaskDINO (Li et al., 2023) R50 52.1M 586 242 46.3 69.0 50.7 26.1 49.3 66.1
@@
Evaluating RF-DETR on RF100-VL. RF100-VL is a challenging detection benchmark composed of 100 diverse datasets. We report latencies, FLOPs, and accuracy averaged over all 100 datasets in Table 4. Our results show that RF-DETR (2x-large) outperforms GroundingDINO and LLMDet while requiring only a fraction of their runtime. Interestingly, RT-DETR outperforms D-FINE (which is built on RT-DETR) at AP50, indicating that D-FINE’s hyperparameters are potentially overfit to COCO. We also note that RF-DETR benefits from scaling to larger backbone sizes (Appendix E). In contrast, YOLOv8 and YOLOv11 consistently underperform DETR-based detectors, and scaling these model families to larger sizes does not improve their performance on RF100-VL (Figure 1).
@@
YOLOv8 (Jocher et al., 2023) N 3.2M 8.7 2.6 55.0 81.1 59.5 4.8 44.1 48.0
YOLOv11 (Jocher et al., 2024) N 2.6M 6.5 3.0 55.5 81.3 60.3 4.7 44.4 49.2
YOLOv8 (Jocher et al., 2023) S 11.2M 28.6 3.1 56.3 82.0 60.9 6.1 45.6 48.6
YOLOv11 (Jocher et al., 2024) S 9.4M 21.5 3.3 56.4 82.5 61.3 6.5 45.5 48.5
YOLOv8 (Jocher et al., 2023) M 25.9M 78.9 5.4 56.5 82.3 60.9 6.4 45.7 48.6
YOLOv11 (Jocher et al., 2024) M 20.1M 68.0 5.1 57.0 82.5 61.9 7.3 46.1 48.6
@@
GroundingDINO (Liu et al., 2023) T 173.0M 1008.3 309.9 62.3 88.8 67.8 39.2 57.7 69.5
LLMDet (Fu et al., 2025) T 173.0M 1008.3 308.4 62.3 88.3 67.8 39.1 57.6 70.3
@@
LW-DETR (Chen et al., 2024a) N 12.1M 21.4 1.9 57.1 84.7 61.5 31.2 51.8 65.8
D-FINE (Peng et al., 2024) N 3.8M 7.3 2.0 58.2 84.4 62.5 32.4 52.9 65.8
RF-DETR (Ours) N 31.2M 34.5 2.5 57.8 85.1 62.5 30.1 52.2 67.2
RF-DETR w/ Fine-Tuning (Ours) N 31.2M 34.5 2.5 58.6 85.7 63.0 31.0 53.2 67.6
LW-DETR (Chen et al., 2024a) S 14.6M 31.8 2.6 57.4 85.0 62.0 32.1 52.1 65.8
D-FINE (Peng et al., 2024) S 10.2M 25.2 3.5 60.3 85.3 65.4 36.6 56.0 68.4
RF-DETR (Ours) S 33.5M 62.4 3.7 60.9 87.5 66.1 34.2 55.7 69.6
RF-DETR w/ Fine-Tuning (Ours) S 33.5M 62.4 3.7 61.2 87.7 66.1 34.9 55.6 69.5
RT-DETR (Zhao et al., 2024) M 36.0M 100.0 4.3 59.6 85.7 64.6 36.4 54.6 67.3
LW-DETR (Chen et al., 2024a) M 28.2M 83.9 4.3 59.8 86.8 64.9 34.0 54.4 68.9
D-FINE (Peng et al., 2024) M 19.2M 56.6 5.6 60.6 85.5 65.8 36.0 56.6 67.5
RF-DETR (Ours) M 33.5M 86.7 4.6 61.7 88.0 66.9 35.8 56.5 70.0
RF-DETR w/ Fine-Tuning (Ours) M 33.5M 86.7 4.6 62.0 88.1 67.1 36.2 56.4 70.2
@@
Table 5: Ablation on Neural Architecture Search. We ablate the impact of each “tunable knob” on accuracy and latency below. Using a gentler set of hyperparameters compared to LW-DETR (e.g. smaller batch size, lower learning rate, replacing batch norm with layer norm) reduces performance by 1%. However, we regain this lost performance by replacing LW-DETR’s CAEV2 backbone with DINOv2. Importantly, the lower learning rate and layer-norm allow us to better preserve DINOv2’s foundational knowledge and allows us to train with larger batch sizes, making weight-sharing NAS more effective. Counterintuitively, introducing weight sharing NAS to the training scheme improves the performance of the base configuration even though patch size 14 isn’t in the NAS search space.
@@
Impact of Backbone Architecture and Pre-Training. We study the impact of different backbone architectures in RF-DETR. We find that DINOv2 achieves the best performance, outperforming CAEv2 by 2%. Interestingly, despite having fewer parameters than SigLIPv2, SAM2’s Hiera-S backbone is considerably slower. This is in contrast with Hiera’s claim that it is meaningfully faster than equivalently performant ViTs. However, Hiera does not explore latency in the context of Flash Attention kernels, which are highly optimized in compilers such as TensorRT. Additionally, existing foundation model families typically do not release lightweight ViT variants such as ViT-S or ViT-T, making it difficult to repurpose such models for real-time applications.
Table 6: Ablation on Backbone. We ablate the impact of using different backbone architectures for RF-DETR below. We find that DINOv2 achieves the highest performance, outperforming CAEv2 by 2.4%. All models are pretrained with 60 epochs of Objects-365 and the “Gentler Hyperparameters” setting. Note that SAM2 and SigLIPv2 perform poorly when evaluated in FP16. Therefore, we report FP16 TensorRT latency with FP32 ONNX accuracy for these two models as an upper bound on their performance if optimized for FP16.
@@
Rethinking Standard Accuracy Benchmarking Practices. Following prior work, we report all COCO results on the validation set. However, relying solely on the validation for both model selection and evaluation can lead to overfitting. For example, D-FINE (which builds on RT-DETR) conducts an extensive hyperparameter sweep on COCO’s validation set and reports its best model. However, evaluating this configuration on RF100-VL shows that D-FINE underperforms RT-DETR on the test set. In contrast, our method achieves state-of-the-art performance among all real-time detectors on both RF100-VL and COCO, demonstrating the robustness of our weight-sharing NAS. In addition to evaluating on COCO, we advocate that future detectors should also evaluate on datasets with public validation and test splits like RF100-VL.
Limitations. Despite controlling for power throttling and GPU overheating during inference, our latency measurements still have a variance of up to 0.1ms due to the non-deterministic behavior of TensorRT during compilation. Specifically, TensorRT can introduce power throttling, which in turn affects the resulting engine and leads to random fluctuations in latency. Although the measurement of a given TensorRT engine is generally consistent, recompiling the same ONNX artifact can produce slightly different latency results. Therefore, we only report latencies with one digit of precision after the decimal place.